OpenStudy (anonymous):

The langth of a rectangle is twice the width. The area is 450yd^2. find the langhth and width.

OpenStudy (vishal_kothari):

let the width by x yd, let the length be 2x yd, Area = w x l = (x)(2x) 450 = 2x^2 225 = x^2 x^2 - 225 = 0 (x-15)(x+15) = 0 therefore, x = -15 (N.A., rejected) , x =15 thus x = 15 Length = 30 yd Width = 15 yd.....

OpenStudy (anonymous):

thank you